OpenStudy (anonymous):

hello anyone good with vectors and finding the resultant? 1.) find the resultant a. 11 units, south, and 6 units west b. 7.5 units north, and 3.5 units east c. 500 units west, and 400 units north d. 7 units east, and 12 units south

OpenStudy (matt101):

You can find the answers to all these questions by sketching out the two vectors you're given, then drawing in the resultant to form a right angled triangle. Since the resultant vector is the hypotenuse of the triangle, you can find its magnitude using the Pythagorean Theorem. Then you'll need to use some trigonometry to find the angle (the direction).

OpenStudy (anonymous):

Aright cool, thank you. my teacher was not the best on explaining how to do it.

OpenStudy (matt101):

Glad I could help. If you want to run your answers by me when you're done I'm happy to check 'em over!

OpenStudy (anonymous):

i do have another question actually. since the Pythagorean theorem gets the hypotenuse. whats the formula for the angle? is it sin or cosin?

OpenStudy (shamim):

U can use Sin Or cosine

OpenStudy (shamim):

Can u tell me the result of ur first wuestion!

OpenStudy (shamim):

U know the formula of pythagorean theorem Hypotenuse^2=leg^2+oposit^2

OpenStudy (matt101):

You can use sin, cos, or tan, as long as you use the right combination of sides in your equation.

OpenStudy (anonymous):

ok so for a.) i got 12.5, 29 degrees, west of south. would that be right?

OpenStudy (shamim):

12.5 is correct!

OpenStudy (shamim):

Tan theta=6/11 Theta=?

OpenStudy (shamim):

I think u can solve these problems by using parallelogram law too

OpenStudy (anonymous):

alrighty this is what i got a) 12.5, 29 degrees west of south b) 8.3, 65 degrees east of north c) 640, 51 degrees north of west d) 14, 60 degrees east of west

OpenStudy (shamim):

I dont hv calculator!

OpenStudy (anonymous):

uh oh haha

OpenStudy (shamim):

I think all r correct

OpenStudy (shamim):

Glad to help !!!!

OpenStudy (anonymous):

yeah thank you

OpenStudy (shamim):

Welcome!!!

OpenStudy (shamim):

Medal????!!!!